Consider a 10-year bond with current price of $98.4 and a duration of 9.2 years. Suppose the yield on the bond is 9.8% per year with continuous compounding. What is the change in the price of the bond if the yield increases by 0.4%?

(required precision: 0.01 +/- 0.01)
Answer #1

We know that change in price = -D*P*change in interest rate

D = -98.4 * 9.2 * 0.004 = $-3.62

So the price will approximately decrease by $3.62 if yield increases by 0.4%